The annual sales for​ Salco, Inc. were $4.49 million last year. The​ firm's end-of-year balance sheet was as​ follows Current assets $507,000 Liabilities $1,002,500 Net fixed assets 1,498,000 Owners' equity 1,002,500 Total Assets $2,005,000 Total $2,005,000

Salco's income statement for the year was as​ follows

Sales $4,490,000

Less: Cost of goods sold (3,491,000)

Gross profit $999,000

Less: Operating expenses (508,000)

Net operating income $491,000

Less: Interest expense (92,000)

Earnings before taxes $399,000

Less: Taxes (35%) (139,650)

Net income $259,350

a. Calculate​ Salco's total asset​ turnover, operating profit​ margin, and operating return on assets. Round to two decimal​ places.

b.  Salco plans to renovate one of its plants and the renovation will require an added investment in plant and equipment of $1.04 million. The firm will maintain its present debt ratio of 50 percent when financing the new investment and expects sales to remain constant. The operating profit margin will rise to 13.4 percent. What will be the new operating return on assets ratio​ (i.e., net operating income divided by÷total ​assets) for Salco after the​ plant's renovation?

c.  Given that the plant renovation in part ​(b​) occurs and​ Salco's interest expense rises by $47,000 per​ year, what will be the return earned on the common​ stockholders' investment? Compare this rate of return with that earned before the renovation. Based on this​ comparison, did the renovation have a favorable effect on the profitability of the​ firm?
Answer #1

a) Total Asset turnover ratio = Net Sales / Total Assets = 4,490,000 / 2,005,000 = 2.24 times

Operating Profit Margin = Operating Profit or Net Operating Income / Net Sales = 491,000 / 4,490,000 = 0.1094 or 10.94%

Operating Return on Assets = Net operating Income / Total Assets = 491,000 / 2,005,000 = 0.2449 or 24.49%

b) Net Fixed Assets will increase by $1.04 million.

So, the year end balance of the Total assets will be = 2,005,000 + 1,040,000 = $3,045,000

The firm will maintain its present debt ratio of 50%

So, the year end balance of Debt (Liabilities) is = 3,045,000*50% = $1,522,500

Owners equity = Total Assets - Liabilities = 3,045,000 - 1,522,500 = $1,522,500

New Operating Profit Margin = 13.4% (Given)

Sales will remain constant at $4,490,000

Net operating income after renovation = Sales*New operating profit margin = 4,490,000*13.4% = $601,660

New Operating Return on asset ratio = Net operating Income / Total Assets = 601,660 / 3,045,000 = 0.1976 or 19.76%

c) Before Renovation:

Net Income = $259,350

Owners' Equity = $1,002,500

Return earned on common stockholders' investment before renovation = Net Income / Owners' Equity = 259,350/1,002,500 = 0.2587 or 25.87%

After Renovation:

Net operating income = 601,660 (Calculated in Part b)

Interest expense rises by $47,000 per year (Given)

New Interest expense = 92,000(before renovation) + 47,000 = $139,000

Earnings before taxes = Net Operating Income - Interest expense = 601,660 - 139,000 = $462,660

Taxes (35%) = 462,660*35% = $161,931

Net Income = Earnings before taxes - Taxes = 462,660-161,931 = $300,729

Owners' Equity = $1,522,500 (Calculated in Part b)

Return earned on common stockholders' investment after renovation = Net Income / Owners' Equity = 300,729/1,522,500 = 0.1975 or 19.75%

The renovation did not have a favourable effect on the profitability of the firm as it reduced the return on common stockholders' investment from 25.87% to 19.75%.